[Brahms-l] Shift report 20050527 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri May 27 2005 - 00:04:09 EDT
Shifter leader was ejk
Shifters were tw


1. Store 6:40pm-7:20pm.
   Beam aborted 7:20pm-8:50pm.
   Last run stopped early for polarization measurement (11:50pm).

2. Useless runs: 14971, 14973, 14974
   Data for 14972, 14975-14980

3. Following BB tubes HV off:
   BBR: tube 31-35
   BBL: tube 40-44

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri May 27 00:04:33 2005

This archive was generated by hypermail 2.1.8 : Fri May 27 2005 - 00:04:48 EDT